### v5

@@ 3.3 Rice_LRT
Our second proposed approach, Rice_LRT, is composed of two steps as illustrated in Fig. 3. The first step involves designing a prompt that is effective for addressing LRT. The second step then applies the prompt to instruct an LLM to predict trace links. We elaborate on these steps next.
@@
In this step, we designed the prompt following recent best practices reported in the RE literature vogelsang2024using ; vogelsang2024specifications . Fig. 4 presents our final prompt, obtained through iterative refinements. The prompt was used on each requirement in the input RD, with the requirement located at the end of the prompt. To design the prompt, we followed the Rice (Role, Instruction, Context, Constraints, Examples) framework, with some small adaptation to the LRT task, as we discuss below. The prompt is structured in the following five elements:
  • Context: This element introduces the LRT task. Since the role is implicitly indicated as a requirements analyst building the trace links, this element subsumes the Role element in the original Rice framework and simply provides the Context. We omitted the explicit mention of the role to obtain a more general applicability of the prompt. For LRT, multiple analysts with different backgrounds are likely involved, e.g., a legal analyst in addition to the requirements analyst. The context corresponds to the text shaded in cyan in Fig. 4.
  • Examples: This element provides a few examples selected from our ground truth. The examples should cover different trace links. Each example is composed of a requirement and the set of trace links alongside the rationale behind each trace link. We note that the LRT task is complex, as we demonstrate throughout the paper. For this reason, we opted for the few-shot prompting technique. This element matches Examples in Rice. Examples corresponds to the text shaded in pink in Fig. 4.
  • Instruction: This element provides explicit instructions on how to perform the LRT task. This element aims to guide the model through the right reasoning process to generate the desired output. The Instruction element corresponds to the text shaded in olive green in Fig. 4. Compared to the original Rice framework, this element contains both the Instruction element combined with the Constraint. The reason for this is that both elements are intertwined in our context. The prompt must therefore account for task-specific considerations, explained below.
@@
      The prompt should encourage the LLM to equally consider other provisions, since only a subset of the provisions are explicitly explained via the examples and rationales in the Examples element. Ideally, the prompt should present an example on each provision. However, this is infeasible since only relevant provisions should be traced to software requirements in a given project. For instance, if the legal basis for collecting personal data is the contract, then unlike explicit consent, only certain data subject rights are applicable according to GDPR and must be appropriately implemented in the software.
@@
      The prompt should account for indirect trace links. As stated above, the LRT is challenging primarily due to the terminology gap between requirements and provisions. We therefore encourage the LLM to use its reasoning capability to identify indirect links, generalizing beyond the provided examples in the prompt.
@@
We note that prompt templates such as Rice serve as illustrative examples or starting points rather than scientifically validated or universally optimal configurations vogelsang2024specifications ; huang2025prompt . Prompt templates are typically task-specific and heuristic in nature. Based on our preliminary analysis of Rice, providing the example before instructions and contextualizing the instructions turned out to work better rather than providing the example at the end.
@@
This step applies the prompt designed in Step 1 to instruct the LLM to predict trace links in textual requirements. A prerequisite for using our prompt involves creating a few examples that will demonstrate the LRT task to the LLM. To effectively trigger the reasoning of the LLM, we built five examples by exposing both the labels (i.e., trace links) and the rationale behind selecting these labels. The examples are then integrated into the above LRT prompt template. We applied our prompt by querying the GPT-4o model through the OpenAI API.
@@ 4.1 Research Questions (RQs)
RQ2. How accurate is $\mathcal{K}$ashif compared to an existing baseline on a standard dataset from the literature? RQ2 aims to assess the value of utilizing ST as enabling technology for addressing the LRT problem compared to a baseline from the existing literature, which we re-implement in this work. The baseline is a classifier that leverages the terminology probability distributions to compute the likelihood that a requirement can be traced to a provision, based on the occurrence of some indicator terms within that provision. The investigation of RQ2 is conducted using the HIPAA dataset.
RQ3. How accurately does $\mathcal{K}$ashif perform on a more complex dataset, spanning multiple requirements types and domains? In RQ3, we test $\mathcal{K}$ashif on four different documents, two shall-requirements and two user stories, covering various domains. These documents are traced to the GDPR privacy requirements. The goal of RQ3 is to investigate the performance of $\mathcal{K}$ashif on a more realistic dataset that captures the complexity of the legal domain.
@@ 4.2 Datasets
We develop our approach and base our initial evaluation on the HIPAA dataset, a publicly available dataset, created and released in 2010 cleland:2010 and reused in 2017 Guo:17 . The dataset was manually created by identifying trace links of requirements against the regulatory statements elicited from the USA government’s Health Insurance Privacy and Portability Act (HIPAA) regulation. The provisions are the following: access control (AC), audit controls (AUD), person or entity authentication (PA), transmission security (TS), unique user identification (UUI), emergency access procedure (EAP), automatic logoff (AL), encryption and decryption (SED), encryption (TED), and integrity controls (IC). HIPAA consists of 10 requirements documents, all shall-requirements, from the healthcare domain. In total, the dataset contains 1,891 requirements, of which 243 have trace links. Table 1 summarizes the different documents (rows) in HIPAA, their description, and the distribution of the trace links across provisions (columns).
@@
To address RQ3, we create and curate four documents covering different requirements types and domains. These documents represent a snapshot of a practical scenario that exemplifies the potential complexity of LRT in practice. For each document, we manually identify trace links between software requirements and a list of 26 provisions derived from GDPR and pertinent to software. Building on existing work Amaral:21 ; Amaral:23a , the codes were comprehensively created, in collaboration with a legal expert (non-author), to represent the privacy requirements in GDPR pertinent to software engineering. Table 2 and 3 describe our test documents. Two co-authors of this paper, with more than 10 years of expertise in requirements engineering, manually analyzed the four documents and identified the trace links for all requirements. The test documents originate from four independent sources, each from a different domain, as shown in Table 2.
It is worth mentioning that the test documents are newly created resources that have not yet been publicly released. Although the underlying content may not rely entirely on new concepts, the traceability links between the software projects and the GDPR are new and have not been shared before. Therefore, we can confidently claim that there is no data leakage when addressing this research question using GPT4.
@@ 4.3 Baselines
To evaluate our proposed solutions, we compare them against carefully selected baselines. we present the baselines for both $\mathcal{K}$ashif and Rice_LRT, covering their classifier-based and prompt-based variants.
@@ 4.3.1 Classifier Baselines
LSI and LDA. Vector Space Model (VSM), LSI, and LDA have been widely adopted in the field of requirements engineering, particularly for addressing the traceability problem guo2025natural , which aims to identify and recover links between related software artifacts such as requirements, design documents, and code. VSM represents each document (e.g., a requirement or a provision) as a vector in a multi-dimensional space, where each dimension corresponds to a term from the overall vocabulary. Similarity between two artifacts is typically computed using the cosine of the angle between their corresponding vectors, capturing the degree of term overlap across all artifacts. LSI extends the Vector Space Model by applying Singular Value Decomposition (SVD) baeza1999modern to the term-document matrix, thereby generating a lower-dimensional latent semantic space. The dimensionality of this subspace is controlled by a manually tuned parameter, commonly referred to as the k value. By capturing the most significant underlying patterns in term usage, LSI implicitly accounts for frequently co-occurring terms, thereby mitigating the vocabulary mismatch problem commonly encountered in information retrieval-based traceability recovery. LDA shares LSI’s goal of uncovering the latent semantic meaning in text, but instead of using SVD, it employs a Bayesian model to identify the underlying structure in word usage across documents.
@@
B. We re-implement as part of this work the baseline B from the literature cleland:2010 ; Guo:17 . B is a probabilistic approach based on occurrences of words in requirements texts and how likely these words are associated with specific provisions. Specifically, B predicts whether a requirement is traced to a provision by identifying keywords (also known as indicator terms) that are present in the requirement. Given an input requirement for which the trace link should be predicted, B requires a training set from which likelihood estimates of indicator terms are computed for the input requirement, representing how likely it is to be relevant to a specific regulation. The training set comprises provisions, software requirements, and the trace links between them. During training, indicator terms are identified and weighted for each provision by parsing the textual requirements traced to these statements. The weights are computed based on factors such as term frequency in related requirements, the fraction of regulation-related requirements containing the term, and the fraction of projects (specific to the HIPAA dataset) involving regulation-related requirements that also contain the term. Given the absence of publicly released implementation for the baseline, we present in this paper a replicated version of B which follows the same procedure described above.
@@
GloVe. We implement a technique based on static word embeddings using GloVe, which generates dense vector representations of words by leveraging global co-occurrence statistics from a large corpus, thereby capturing both semantic and syntactic relationships between words. In our work, we use GloVe as a trainable embedding layer and as the input layer in a simple neural network architecture (with a single layer to predict the score). In this context, the inputs consist of a requirement and a candidate provision, and the output is a probability score computed using a sigmoid activation function. The model is subsequently fine-tuned on the training dataset, allowing it to adapt the pre-trained word embeddings to the specific characteristics of the traceability task.
@@
TraceBERT. we utilize the publicly available implementation of TraceBERT, a transformer-based model that, like our approach, is built on a transformer architecture. In this paper, the authors propose three distinct BERT-based architectures (TWIN, Single, and Siamese) for fine-tuning the pre-trained BERT model specifically for the traceability link prediction task. Each architecture represents a different way to structure the input and optimize the model for semantic similarity between software artifacts. Based on reported evaluation results, the Single-TraceBERT architecture outperforms the other variations, and we adopt it as one of the baseline models in this work. Note that, unlike the original TraceBERT paper, which includes intermediate training on CodeSearchNet husain2019codesearchnet , we skipped this step as it is not relevant to legal domain traceability as CodeSearchNet contains programming language artifacts which is beneficial for source code and issue traceability, but differs fundamentally—in both terminology and syntactic structure—from establishing traceability between two NL artifacts in our case. Therefore, we directly fine-tuned the authors’ architecture on our dataset.
@@
LLaMa. Finally, we utilize the publicly available implementation of a recent approach by Ge et al. ge2025cross , which primarily focused on establishing traceability between low- and high-level requirements using LLaMa2 touvron2023llama . In this work, they use data augmentation to fine-tune LLaMa2 at different scales (1.1B, 7B, and 13B) and demonstrate that LLaMa2-7B achieves the best performance. They demonstrate that the fine-tuned LLaMa outperforms both IR-based and ML-based approaches on six datasets, using a 90/10 train–test split. In our study, we evaluate how well this approach performs on the requirement–provision task under a more complex evaluation setting.
@@ 4.3.2 Prompt Baselines
For Rice_LRT, we incorporate four baselines, P1, P2, P3_1, and P3_2, from the literature hey2025requirements ; ronanki2024requirements , which focus on identifying links between high-level and low-level requirements. This comparison enables us to assess the effectiveness of our context-rich prompt, Rice_LRT, relative to recent prompt-based techniques reported in the literature for traceability.
@@
P3. We adopt two prompt templates used by Ge et al. ge2025cross as part of their investigation into high- and low-level requirement traceability. Their study evaluated four zero-shot prompt variations, ranging from a simple prompt to a more detailed CoT formulation. In order to have a variety of prompts in our baselines, we employ the prompt with the best results in their work, we call it P3_1 throughout the paper, along with the CoT-based prompt as it includes more detailed instructions compared to P1 and P2, we call it P3_2. We modify the prompt to better fit the LRT task. The prompts we used are as follows:
P3-1: Consider the following scenario where the Requirement text represent requirements of the software system and Regulation represent a GDPR regulation. Let’s think step by step: 1. Identify the key elements of the requirement; 2. Match these elements with the regulation; 3. Highlight any gaps; 4. Infer implicit traceability; Finally, answer with Yes or No if there is a traceability relationship, and provide a brief rationale in about 10 words.
@@
P3-2: Consider the following scenario where the Requirement text represent requirements of the software system and Regulation represent a GDPR regulation. Answer with Yes or No: Does the regulation trace back to the requirement?
@@ 4.4 Implementation
$\mathcal{K}$ashif. We implement $\mathcal{K}$ashif in Python 3.8. For preprocessing the text, we use the NLTK toolkit (v 3.8.1). We access the ST pre-trained models through the Hugging Face Transformers library (4.44.0). For fine-tuning, we use the Sentence-Transformers library (2.6.1). We use the same library for computing cosine similarity. Our experiments were performed on an RTX 6000 GPU with 24 GB of RAM. For fine-tuning, we relied on the ST29 base model. We performed grid search to tune the hyperparameters, considering epochs=[5, 10, 15], batch size=[8, 16, 32], optimizer=[AdamW, RMSprop], learning rate=[2e-5, 2e-3, 2e-2], weight decay=[2e-3, 2e-2, 2e-1], and warmup percentage=[5e-2, 5e-1]. The best-performing hyperparameters were epochs=5, batch size=8, optimizer=AdamW, linear learning rate=2e-5, weight decay=2e-1, and warmup percentage=5e-2. Checkpoints were saved at the end of each epoch, and the final model was selected based on the lowest observed loss.
@@
LSI and LDA. For LSI and LDA, we utilized the scikit-learn library (v1.7.0), setting the number of components to 50 within a search space of 5, 10, 25, and 50 components. For LDA the Document–Topic Dirichlet Prior ($\alpha$) and Topic–Word Dirichlet Prior ($\beta$) had the values of 2e-2 within a search space of 1e-2, 2e-2, and 1e-1. For both models, we employed TF-IDF as the feature weighting scheme using scikit-learn library (v1.7.0). Our experiments showed that the LSI model achieved optimal performance when stopwords and punctuation were removed. In contrast, the LDA model performed best when stopwords and punctuation were removed and lemmatization was applied using NLTK toolkit (v3.8.1). We did not apply minimum cutoff thresholds for the vocabulary, as filtering out low-frequency terms could risk omitting important information relevant to some regulatory provisions. For the maximum document frequency cutoff, we experimented with values of 0, 1e-1, and 1. We observed that as the cutoff value increased, the model’s performance decreased. Therefore, we decided not to apply a maximum cutoff in our configuration.
@@
TraceBERT. For TraceBERT, we used the publicly available code from data preparation for tuning the model and retained the fine-tuning settings in its shared implementation.
RoBERTa. For RoBERTa, the entire fine-tuning setup is implemented using the Hugging Face Transformers library. We performed grid search to tune the hyperparameter, considering using epochs=[5, 10, 15], batch size=[8, 16, 32], optimizer=[AdamW, RMSprop], learning rate=[1e-5, 2e-5, 1e-3, 2e-3], and weight decay=[0, 1e-5, 1e-3, 1e-2, 1e-1]. The best-performing hyperparameters were epochs=5, batch size=32, optimizer=AdamW, learning rate=2e-5, and weight decay=1e-2.
LLaMa. For LLaMa, we used the publicly available implementation presented by Ge et al ge2025cross . We relied on their released code to generate the training set following the procedure outlined in their paper. Since using the same hyperparameters resulted in a single label being produced for all pairs, we performed another round of hyper-parameter tuning considering learning rate=[2e-8, 2e-7, 2e-5, 2e-5, 2e-4, 2e-3, 2e-2] and weight decay=[0, 1e-5, 1e-3, 1e-2, 1e-1], which the best-performing hyperparameters were learning rate=2e-4 and weight decay=1e-5.
P1. For P1, we re-implemented the approach following the settings provided by the authors. The prompt was executed on GPT-4o with a temperature of 0 and a random seed of 133742243. For the RAG step, we used the same embeddings as the authors proposed, specifically the OpenAI embedding text-embedding-3-large.
P2. For P2, we re-implemented the approach following the settings provided by the authors. While the original paper used GPT-3.5-turbo, our comparative experiments revealed that GPT-4o yielded better outcomes. Therefore, the results for P2 presented in this paper are based the GPT-4o model. Moreover, the choice of random seed and temperature followed the best-performing settings reported in their paper.
P3. For P3_1 and P3_2, we re-implemented the approach using the settings provided by the authors over GPT-4o.

### v6

@@ 3.3 Rice_LRT
Our second proposed approach, Rice_LRT, comprises two steps, as illustrated in Fig. 3. The first step is to design a prompt effective for addressing LRT. The second step then applies the prompt to instruct an LLM to predict trace links. We elaborate on these steps next.
@@
In this step, we designed the prompt in accordance with recent best practices reported in the RE literature vogelsang2024using ; vogelsang2024specifications . Fig. 4 presents our final prompt, obtained through iterative refinements. The prompt was applied to each requirement in the input RD, with the requirement listed at the end of the prompt. To design the prompt, we followed the Rice (Role, Instruction, Context, Constraints, Examples) framework, with some small adaptations to the LRT task, as we discuss below. The prompt is structured in the following five elements:
  • Context: This element introduces the LRT task. Since the role is implicitly indicated as a requirements analyst building the trace links, this element subsumes the Role element in the original Rice framework and simply provides the Context. We omitted the explicit mention of the role to obtain a more general applicability of the prompt. For LRT, multiple analysts with diverse backgrounds are likely involved, for example, a legal analyst in addition to a requirements analyst. The context corresponds to the text shaded in cyan in Fig. 4.
  • Examples: This element provides a few examples selected from our ground truth. The examples should cover trace links to different provisions. Each example comprises a requirement and the set of trace links, along with the rationale for each trace link. We note that the LRT task is complex, as we demonstrate throughout the paper. For this reason, we opted for the few-shot prompting technique. This element matches Examples in Rice. Examples corresponds to the text shaded in pink in Fig. 4. The few-shot examples are not present in the test data.
  • Instruction: This element provides explicit instructions on how to perform the LRT task. This element aims to guide the model through the right reasoning process to generate the desired output. The Instruction element corresponds to the text shaded in olive green in Fig. 4. Compared with the original Rice framework, this element combines the Instruction and Constraint elements. This is because the two elements are intertwined in our context. The prompt must therefore account for task-specific considerations, explained below.
@@
      The prompt should encourage the LLM to equally consider other provisions, since only a subset of the provisions are explicitly explained via the examples and rationales in the Examples element. Ideally, the prompt should present an example of each provision. However, this is infeasible, as only the relevant provisions should be traced to the software requirements for a given project. For instance, if the legal basis for collecting personal data is the contract, then, unlike explicit consent, only certain data subject rights apply under the GDPR and must be appropriately implemented in the software.
@@
      The prompt should account for indirect trace links. As noted above, the LRT is challenging primarily due to the terminology gap between the requirements and the provisions. We therefore encourage the LLM to use its reasoning capabilities to identify indirect links and to generalize beyond the examples provided in the prompt.
@@
We note that prompt templates such as Rice serve as illustrative examples or starting points rather than scientifically validated or universally optimal configurations vogelsang2024specifications ; huang2025prompt . Therefore, during the refinement process, we randomly selected a separate subset from the few-shot example pool (two out of five examples per document; for a total of eight examples) and used them as inputs for refinement, without including them inside the prompt. In each iteration, a requirement text from this subset was used to obtain the model’s response, and the generated rationale was then analyzed to refine the instructions by adjusting constraints, ensuring that the resulting rationales were non-hallucinatory, less ambiguous, and judged as logical by manual evaluation. Note that, after refinement, these examples were returned to the few-shot sets and used as few-shot examples during the test stage. Moreover, the analysis indicates that presenting the example before the instructions and contextualizing the instructions were more effective than presenting the example at the end.
@@
This step applies the prompt designed in Step 1 to instruct the LLM to predict trace links in textual requirements. A prerequisite for using our prompt involves creating a few examples that will demonstrate the LRT task to the LLM. To effectively trigger the reasoning of the LLM, we built five examples by exposing both the labels (i.e., trace links) and the rationale for selecting them. The examples are then integrated into the above LRT prompt template. We applied our prompt by querying the GPT-4o model through the OpenAI API.
@@ 4.1 Research Questions (RQs)
RQ2. How accurate is $\mathcal{K}$ashif compared to an existing baseline on a standard dataset from the literature? RQ2 aims to assess the value of using ST as an enabling technology for addressing the LRT problem, compared with a baseline from the existing literature that we reimplement in this work. The baseline is a classifier that leverages the terminology probability distributions to compute the likelihood that a requirement can be traced to a provision, based on the occurrence of some indicator terms within that provision. The investigation of RQ2 is conducted using the HIPAA dataset.
RQ3. How accurately does $\mathcal{K}$ashif perform on a more complex dataset, spanning multiple requirements types and domains? In RQ3, we test $\mathcal{K}$ashif on four different documents, two shall-requirements and two user stories, covering various domains. These documents are derived from the GDPR’s privacy requirements. The goal of RQ3 is to evaluate kashif’s performance on a more realistic dataset that captures the complexity of the legal domain.
@@ 4.2 Datasets
We develop our approach and base our initial evaluation on the HIPAA dataset, a publicly available dataset, created and released in 2010 cleland:2010 and reused in 2017 Guo:17 . The dataset was manually created by identifying trace links of requirements against the regulatory statements elicited from the USA government’s Health Insurance Privacy and Portability Act (HIPAA) regulation. The provisions are the following: access control (AC), audit controls (AUD), person or entity authentication (PA), transmission security (TS), unique user identification (UUI), emergency access procedure (EAP), automatic logoff (AL), encryption and decryption (SED), encryption (TED), and integrity controls (IC). HIPAA comprises 10 requirements documents, all of which are shall requirements, from the healthcare domain. In total, the dataset contains 1,891 requirements, of which 243 have trace links. Table 1 summarizes the different documents (rows) in HIPAA, their description, and the distribution of the trace links across provisions (columns).
@@
To address RQ3 and RQ4, we create and curate four documents covering different requirements types and domains. These documents provide a snapshot of a practical scenario that exemplifies the potential complexity of LRT. Moreover, we limit testing of the prompts and LLMs to unseen datasets to minimize the risk of data leakage. For each document, we manually identify trace links between software requirements and a list of 26 provisions derived from GDPR and pertinent to software. Building on existing work Amaral:21 ; Amaral:23a , the codes were comprehensively created, in collaboration with a legal expert (non-author), to represent the privacy requirements in GDPR pertinent to software engineering. Table 2 and 3 describe our test documents. Two co-authors of this paper, with more than 10 years of expertise in requirements engineering, manually analyzed the four documents and identified the trace links for all requirements. The test documents originate from four independent sources, each from a different domain, as shown in Table 2.
It is worth mentioning that the test documents, unlike HIPAA dataset, are newly created resources that have not yet been publicly released. Although the underlying content may not rely entirely on new concepts, the traceability links between the software projects and the GDPR are new and have not been shared before. Therefore, we can confidently claim that there is no data leakage when addressing this research question using GPT4.
@@ 4.3 Baselines
To evaluate our proposed solutions, we compare them against carefully selected baselines. We present the baselines for both $\mathcal{K}$ashif and Rice_LRT, including their classifier- and prompt-based variants.
@@ 4.3.1 Classifier Baselines
LSI and LDA. Vector Space Model (VSM), LSI, and LDA have been widely adopted in the field of requirements engineering, particularly for addressing the traceability problem guo2025natural , which aims to identify and recover links between related software artifacts such as requirements, design documents, and code. VSM represents each document (e.g., a requirement or a provision) as a vector in a multi-dimensional space, where each dimension corresponds to a term from the overall vocabulary. Similarity between two artifacts is typically computed using the cosine of the angle between their corresponding vectors, capturing the degree of term overlap across all artifacts. LSI extends the Vector Space Model by applying Singular Value Decomposition (SVD) baeza1999modern to the term-document matrix, thereby generating a lower-dimensional latent semantic space. The dimensionality of this subspace is controlled by a manually tuned parameter, commonly referred to as k. By capturing the most significant underlying patterns in term usage, LSI implicitly accounts for frequently co-occurring terms, thereby mitigating the vocabulary mismatch problem commonly encountered in information retrieval-based traceability recovery. LDA shares LSI’s goal of uncovering the latent semantic meaning in text, but instead of using SVD, it employs a Bayesian model to identify the underlying structure in word usage across documents.
@@
B. We re-implement as part of this work the baseline B from the literature cleland:2010 ; Guo:17 . B is a probabilistic approach based on the occurrences of words in requirements texts and on how likely these words are to be associated with specific provisions. Specifically, B predicts whether a requirement is traced to a provision by identifying keywords (also known as indicator terms) that are present in the requirement. Given an input requirement for which the trace link should be predicted, B requires a training set from which likelihood estimates of indicator terms are computed for the input requirement, representing how likely it is to be relevant to a specific regulation. The training set comprises provisions and software requirements, along with the trace links between them. During training, indicator terms are identified and weighted for each provision by parsing the textual requirements traced to these statements. The weights are computed based on factors such as term frequency in related requirements, the fraction of regulation-related requirements containing the term, and the fraction of projects (specific to the HIPAA dataset) involving regulation-related requirements that also contain the term. Given the absence of publicly released implementation for the baseline, we present in this paper a replicated version of B which follows the same procedure described above.
@@
GloVe. We implement a technique based on static word embeddings using GloVe, which generates dense vector representations of words by leveraging global co-occurrence statistics from a large corpus, thereby capturing both semantic and syntactic relationships between words. In our work, we use GloVe as both a trainable embedding layer and the input layer in a simple neural network architecture (with a single-layer architecture to predict the score). In this context, the inputs comprise a requirement and a candidate provision, and the output is a probability score computed via a sigmoid activation function. The model is subsequently fine-tuned on the training dataset, allowing it to adapt the pre-trained word embeddings to the specific characteristics of the traceability task.
@@
TraceBERT. we utilize the publicly available implementation of TraceBERT, a transformer-based model that, like our approach, is built on a transformer architecture. In this paper, the authors propose three distinct BERT-based architectures (TWIN, Single, and Siamese) for fine-tuning the pre-trained BERT model for the traceability link-prediction task. Each architecture represents a distinct approach to structuring the input and optimizing the model for semantic similarity between software artifacts. Based on reported evaluation results, the Single-TraceBERT architecture outperforms the other variations, and we adopt it as one of the baseline models in this work. Note that, unlike the original TraceBERT paper, which includes intermediate training on CodeSearchNet husain2019codesearchnet , we skipped this step as it is not relevant to legal domain traceability as CodeSearchNet contains programming language artifacts which is beneficial for source code and issue traceability, but differs fundamentally—in both terminology and syntactic structure—from establishing traceability between two NL artifacts in our case. Therefore, we directly fine-tuned the authors’ architecture on our dataset.
@@
LLaMa. Finally, we utilize the publicly available implementation of a recent approach by Ge et al. ge2025cross , which primarily focused on establishing traceability between low- and high-level requirements using LLaMa2 touvron2023llama . In this work, they use data augmentation to fine-tune LLaMa2 at different scales (1.1B, 7B, and 13B) and demonstrate that LLaMa2-7B achieves the best performance. They demonstrate that the fine-tuned LLaMa outperforms both IR- and ML-based approaches across six datasets using a 90/10 train–test split. In our study, we evaluate how well this approach performs on the requirement–provision task under a more complex evaluation setting.
@@ 4.3.2 Prompt Baselines
For Rice_LRT, we incorporate four baselines, P1, P2, P3_1, and P3_2, from the literature hey2025requirements ; ronanki2024requirements , which focus on identifying links between high-level and low-level requirements. This comparison enables us to assess the effectiveness of our context-rich prompt, Rice_LRT, relative to recent traceability prompt-based techniques reported in the literature.
@@
P3. We adopt two prompt templates used by Ge et al. ge2025cross as part of their investigation into high- and low-level requirement traceability. Their study evaluated four zero-shot prompt variations, ranging from a simple prompt to a more detailed CoT formulation. In order to have a variety of prompts in our baselines, we employ the prompt with the best results in their work, which we call P3_1 throughout the paper, along with the CoT-based prompt, as it includes more detailed instructions compared to P1 and P2, which we call P3_2. We modify the prompt to better fit the LRT task. The prompts we used are as follows:
P3-1: Consider the following scenario where the Requirement text represents requirements of the software system and Regulation represents a GDPR regulation. Let’s think step by step: 1. Identify the key elements of the requirement; 2. Match these elements with the regulation; 3. Highlight any gaps; 4. Infer implicit traceability; Finally, answer with Yes or No if there is a traceability relationship, and provide a brief rationale in about 10 words.
@@
P3-2: Consider the following scenario where the Requirement text represents requirements of the software system and Regulation represents a GDPR regulation. Answer with Yes or No: Does the regulation trace back to the requirement?
@@ 4.4 Implementation
$\mathcal{K}$ashif. We implement $\mathcal{K}$ashif in Python 3.8. For text preprocessing, we use the NLTK toolkit (v. 3.8.1). We access the ST pre-trained models through the Hugging Face Transformers library (4.44.0). For fine-tuning, we use the Sentence-Transformers library (2.6.1). We use the same library for computing cosine similarity. Our experiments were performed on an RTX 6000 GPU with 24 GB of RAM. For fine-tuning, we relied on the ST29 base model. We performed grid search to tune the hyperparameters, considering epochs=[5, 10, 15], batch size=[8, 16, 32], optimizer=[AdamW, RMSprop], learning rate=[2e-5, 2e-3, 2e-2], weight decay=[2e-3, 2e-2, 2e-1], and warmup percentage=[5e-2, 5e-1]. The best-performing hyperparameters were epochs=5, batch size=8, optimizer=AdamW, linear learning rate=2e-5, weight decay=2e-1, and warmup percentage=5e-2. Checkpoints were saved at the end of each epoch, and the final model was selected based on the lowest observed loss.
@@
LSI and LDA. For LSI and LDA, we utilized the scikit-learn library (v1.7.0), setting the number of components $(n_{\text{components}}$) to 50 within a search space of 5, 10, 25, and 50 components. For LDA, the Document–Topic Dirichlet Prior ($\alpha$) and Topic–Word Dirichlet Prior ($\beta$) had the values of 2e-2 within a search space of 1e-2, 2e-2, 1e-1, and 2e-1. The ranges for $\alpha$ and $\beta$ were chosen to encompass commonly used settings in the literature uto2017diverse ; silva2021topic , with particular emphasis on values proportional to $1/n_{\text{components}}$. For both models, we employed TF-IDF as the feature weighting scheme using the scikit-learn library (v1.7.0). Our experiments showed that the LSI model achieved optimal performance when stopwords and punctuation were removed. In contrast, the LDA model performed best when stopwords and punctuation were removed, and lemmatization was applied using the NLTK toolkit (v3.8.1). We did not apply minimum cutoff thresholds for vocabulary, as filtering out low-frequency terms could risk omitting information relevant to certain regulatory provisions. For the maximum document frequency cutoff, we experimented with values of 0, 1e-1, and 1. We observed that as the cutoff value increased, the model’s performance decreased. Therefore, we decided not to apply a maximum cutoff in our configuration.
@@
TraceBERT. For TraceBERT, we used the publicly available data-preparation code to tune the model and retained the fine-tuning settings from the shared implementation.
RoBERTa. For RoBERTa, the entire fine-tuning setup is implemented using the Hugging Face Transformers library. We performed a grid search to tune the hyperparameters, considering using epochs=[5, 10, 15], batch size=[8, 16, 32], optimizer=[AdamW, RMSprop], learning rate=[1e-5, 2e-5, 1e-3, 2e-3], and weight decay=[0, 1e-5, 1e-3, 1e-2, 1e-1]. The best-performing hyperparameters were epochs=5, batch size=32, optimizer=AdamW, learning rate=2e-5, and weight decay=1e-2.
LLaMa. For LLaMa, we used the publicly available implementation presented by Ge et al ge2025cross . We relied on their released code to generate the training set following the procedure outlined in their paper. Since using the same hyperparameters resulted in a single label being produced for all pairs, we performed another round of hyperparameter tuning, considering learning rate=[2e-8, 2e-7, 2e-5, 2e-5, 2e-4, 2e-3, 2e-2] and weight decay=[0, 1e-5, 1e-3, 1e-2, 1e-1], where the best-performing hyperparameters were learning rate=2e-4 and weight decay=1e-5.
P1. For P1, we reimplemented the approach using the authors’ settings. The prompt was executed on GPT-4o with a temperature of 0 and a random seed of 133742243. For the RAG step, we used the same embeddings as the authors proposed, specifically the OpenAI embedding text-embedding-3-large.
P2. For P2, we reimplemented the approach using the authors’ settings. While the original paper used GPT-3.5-turbo, our comparative experiments revealed that GPT-4o yielded better outcomes. Therefore, the results for P2 presented in this paper are based on the GPT-4o model. Moreover, the choice of random seed and temperature was based on the best-performing settings reported in their paper.
P3. For P3_1 and P3_2, we reimplemented the approach using the authors’ settings on GPT-4o.